# Relativistic Guiding-Center Motion: Action Principle, Kinetic Theory, and Hydrodynamics

### v1

@@ Abstract
We treat the guiding-center dynamics in a non-uniform external Maxwell field using a manifestly Lorentz covariant action principle which easily reproduces the known Vandervoort equations of motion. We derive the corresponding kinetic theory and ideal hydrodynamic theory. In contrast to conventional five-equation hydrodynamics, the guiding-center hydrodynamics needs only three equations due to a constraint on the motion across magnetic field. We argue that this hydrodynamics applies more generally than the kinetic theory, e.g., for strongly-coupled quark-gluon plasma.
Introduction.—About fifty years ago, Vitaly Ginzburg identified matter in ultrahigh magnetic fields as one of the most important and interesting problems in physics [1]. This problem remains current to this day. In astrophysics, magnetic fields exceeding $10^{15}~{}\text{G}$ have been found on the surface of a type of neutron stars called magnetars [2, 3]. In peripheral ultrarelativistic heavy-ion collisions, the magnetic field created by the colliding nuclei may exceed $10^{19}~{}\text{G}$ [4, 5] and leave measurable effects on the collisions [6]. In condensed matter physics, magnetic fields available in the lab ($\sim 10^{5}~{}\text{G}$) can already be considered strong at low densities of charge carriers and temperatures, and lead to nontrivial phenomena like the integer and fractional quantum Hall effects [7, 8].
It is known that properties of matter change dramatically in high magnetic fields [9]. Ordinary atoms become deformed when the magnetic field $B$ reaches $m_{e}^{2}ce^{3}/\hbar^{3}\approx 2.35\times 10^{9}~{}\text{G}$ and, when $B$ exceeds $m_{e}^{2}c^{3}/(e\hbar)\approx 4.41\times 10^{13}~{}\text{G}$, the cyclotron motion of electrons is relativistic. Electron motion at high $B$ can be decomposed into a fast cyclotron motion and a slower motion of the orbit’s guiding center. At very high $B$, the cyclotron motion is quantum and, in the most extreme case, frozen onto the lowest Landau level. The motion of the guiding center, on the other hand, is (semi-)classical and is a text-book problem of plasma physics. It is well-known that the guiding center moves mostly along the magnetic field line with a slow “drift” off that line when the field is inhomogeneous [10]. This drift motion is important in various physical problems including cosmic-ray transport [11] and thermonuclear plasma confinement [12]. A plasma of electrons in a high $B$ field can be thought of as a plasma of guiding centers, and finding a description of the collective dynamics of these guiding centers is crucial for understanding properties of matter in very high magnetic fields.
@@
Relativistic motion of the guiding center.—In the relativistic setting, the problem of guiding-center motion has been fully solved since the 1960s [13, 14]. Here we provide an alternative, simpler, and manifestly Lorentz covariant description of the motion based on an action principle. This formulation will turn out to be particularly apt for our subsequent derivation of the kinetic and hydrodynamic theories.
We assume the presence of a strong electromagnetic field $F_{\mu\nu}$ which is magnetically dominated, i.e., $F_{\mu\nu}F^{\mu\nu}=2(\bm{B}^{2}-\bm{E}^{2})>0$. At any given spacetime point there exist a frame where $\bm{E}\parallel\bm{B}$. We denote by $B_{*}$ and ${E_{*}}$ the magnitudes of the electric and magnetic fields in that frame (${E_{*}}<0$ if the fields are antiparallel). In terms of the Lorentz invariants, $F_{\mu\nu}F^{\mu\nu}=2(B_{*}^{2}-E_{*}^{2})$, $F_{\mu\nu}\tilde{F}^{\mu\nu}=-4E_{*}B_{*}$  [^1] We use the convention of Refs. [10, 30]: the metric signature is (+,,,)(+,-,-,-)( + , - , - , - ), $A^{\mu}=(\phi,\mathbf{A})$, $\tilde{F}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\alpha\beta}F_{\alpha\beta}$, where $\epsilon^{\mu\nu\alpha\beta}$ is the Levi-Civita tensor, $\epsilon^{0123}=+1$..
We consider the motion of an electron in this field and, to simplify further order-of-magnitude estimates, assume that the motion is relativistic: $v\sim 1$. Let $p_{\perp}$ be the magnitude of the electron’s momentum in the plane perpendicular the magnetic field. Then cyclotron radius is $a=p_{\perp}/(eB_{*})$. We assume $a$ is much smaller than the scale $L$ on which the magnetic field varies: $a\ll L$. In that case the motion of the electron is a superposition of a fast cyclotron motion and the smoother motion of the guiding center.
@@
$B_{\mu\nu}=\frac{F_{\mu\nu}-\epsilon_{*}\tilde{F}_{\mu\nu}}{1+\epsilon_{*}^{2}};\quad E_{\mu\nu}=\epsilon_{*}\frac{\tilde{F}_{\mu\nu}+\epsilon_{*}F_{\mu\nu}}{1+\epsilon_{*}^{2}},$ (1)
@@
The equations governing the motion of the guiding center was first derived in Ref. [13] (see also Ref. [14]) in explicit component form where Lorentz invariance is hidden. We show that these equations can be obtained straightforwardly from the manifestly Lorentz invariant least action principle for a relativistic particle with an additional constraint. Namely, let the trajectory of the particle in phase space be $(x^{\mu}(\tau),p_{\mu}(\tau))$, where $\tau$ is an affine parameter along the worldline. Consider the following Lagrangian,
$L=-\dot{x}\cdot(p+A)+\frac{\alpha}{2}(p^{2}-\tilde{m}^{2})+\lambda_{\mu}B^{\mu\nu}p_{\nu}.$ (2)
@@
together with two constraints $p^{2}=\tilde{m}^{2}$ and $B^{\mu\nu}p_{\nu}=0$. From Eq. (5a) one sees that $\alpha$ and $\lambda_{\alpha}$ depend on the choice oe the affine parameter $\tau$. For example, if $B_{\mu\nu}$ is purely spatial, choosing $\tau=t$ means that $\alpha=1/p_{0}$.
While the momentum of the guiding center always points in the direction of the magnetic field due to the constraint $B^{\mu\nu}p_{\nu}=0$, the particle does not move strictly along the magnetic field lines: it also slowly “drifts” in an orthogonal direction with a velocity given by the spatial components of $v_{\mathrm{D}}^{\mu}\equiv\lambda_{\alpha}B^{\alpha\mu}\,$. Substituting Eq. (5a) into Eq. (5b) and multiplying by $B^{\lambda\mu}$ we find
@@
where we dropped terms higher order in gradients of the magnetic field.
Equation (6) compactly and covariantly summarizes the results of Refs. [13, 14]. The first term in Eq. (6) reproduces the curvature drift. The second one reproduces the gradient drift in time-independent, spatially varying fields, while also providing additional effects proportional to the time derivative of the magnetic field [17]. Since $v_{\mathrm{D}}$, and thus $\lambda$, is proportional to the gradient of $B$, the last term in Eq. (5b) is second order in gradients.
Kinetic equation.—Using the equation of motion for a single particle derived above, we now study the kinetic theory of a collection of charged particles moving in a large external, almost crossed ($\epsilon_{*}\ll 1$), field. To be able to replace the particle motion by that of the guiding centers, the cyclotron radius must be shorter than the mean free path. For example, for a relativistic electron-positron plasma with $T\gtrsim m_{e}$, the cyclotron radius is $T/(eB)$ while the mean free path is (up to a log) $1/(e^{4}T)$. The kinetic theory we are going to derive works when $B\gg e^{3}T^{2}$.
@@
where the Liouville operator $\mathcal{L}[f]$ represents the rate of change (per unit worldline “time”) of the occupation number $f(x,p)$ along the particle worldline determined by Eqs. (5). The occupation number is changed by collisions, therefore the rate is equal to the collision integral $\mathcal{C}[f]$.
To describe transport in the gas we need to know also the invariant phase-space integration measure, i.e., the number of single-particle states in a given phase-space volume at a given time in the lab frame. The number of states is time-independent if the phase-space volume evolves according to equations of motion. This statement can be expressed in Lorentz covariant form by introducing the conserved phase-space 8-current $J^{M}$ given by [18]
@@
i.e., by the product of the phase-space “velocity” and the “density” $W$  [^3] We need to extend the velocities $\dot{x}^{\mu}$, $\dot{p}_{\nu}$ to the whole 8-dimensional phase space. We can do that by simply declaring Eqs. (5) and (6) to be valid everywhere, even outside the constraint hypersurface.
@@
$W=\frac{B_{*}^{3}+B_{*}p_{\mu}\partial_{\nu}B^{\mu\nu}}{2\pi^{2}\alpha}\theta(p_{0})\delta(p^{2}-\tilde{m}^{2})\delta^{2}(B^{\mu\nu}p_{\nu}).$ (10)
Equation $\partial_{M}J^{M}=0$ determines $W$ only up to an overall constant. This constant is chosen in Eq. (10) to match the Landau level degeneracy.
Hydrodynamics.—We can now use the guiding-center kinetic theory to describe the transport in the hydrodynamic regime. For simplicity, we consider a single-component plasma, neutralized by an inert background charge density. The electromagnetic field is treated mostly as a fixed background, but can be made dynamical by supplementing the hydrodynamic equations with Maxwell’s equations.
@@
where we defined the projector $\tilde{\Delta}^{\lambda}_{\nu}$ as
For a magnetic field along the $z$ axis, $\Delta^{\mu}_{\nu}$ projects onto the $(x,y)$ plane, and $\tilde{\Delta}^{\mu}_{\nu}$ onto the $(t,z)$ hyperplane. The presence of $\tilde{\Delta}$ means that Eq. (11b) represents only two independent equations, not four. These conservation laws are supplemented by the constitutive relations that express current $\tilde{N}^{\mu}$ and stress tensor $T^{\mu\nu}$ in terms of three independent hydrodynamic variables: the local temperature $T(x)$, local chemical potential $\mu(x)$, and local velocity $u^{\mu}(x)$. The velocity is subject to the constraints $u\cdot u=1$, $B_{\mu\nu}(x)u^{\nu}(x)=0$ and hence represents only one degree of freedom.
@@
Derivation of hydrodynamics.—Multiplying Eq. (7) by $W$ and $Wp_{\nu}$ and integrating over $p$, using the fact that the collision term $\mathcal{C}[f]$ conserves the particle number (charge) and energy-momentum, i.e.,  [^4] We have chosen the charge of particles to be 1. If there is more than one types of particles, one needs to sum over the particle types $a$. For example, conservation of charge becomes $\sum_{a}\int_{p}Wq_{a}C[f_{a}]=0$ where $q_{a}$ is the charge of the particles of type $a$. Similarly, if the values of adiabatic invariant $J$ vary, they have to be summed over too, e.g., in the second equation of (19).
@@
$F_{\nu}=F_{\nu\lambda}N^{\lambda}-M\partial_{\nu}B_{*},~{}~{}M\equiv-\!\int_{p}W\frac{\alpha J}{2}f.$ (19)
@@
The current in Eq. (17) consists of two parts: $N^{\mu}=\tilde{N}^{\mu}+N_{\text{D}}^{\mu}$, where the longitudinal current
$\tilde{N}^{\mu}\equiv\int_{p}W\alpha p^{\mu}f\,,$ (20)
flows along the magnetic field lines ($B_{\mu\nu}\tilde{N}^{\nu}=0$), while the “drift” current
$N_{\text{D}}^{\mu}\equiv\int_{p}Wv_{\text{D}}^{\mu}f=B_{*}^{-2}(\tilde{T}^{\lambda}_{\nu}\partial_{\lambda}B^{\mu\nu}-MB^{\mu\nu}\partial_{\nu}B_{*})\,,$ (21)
@@
Then Eq. (22) takes a simper form of Eq. (11b). Physically, the contribution $T_{\perp}^{\mu\nu}$ in Eq. (23) comes from the fast transverse motion of the particles, absent in Eq. (18). Indeed,
@@
While $\overline{p_{\perp}^{\mu}}=0$, $\overline{p_{\perp}^{\mu}p_{\perp}^{\nu}}=-\Delta^{\mu\nu}p_{\perp}^{2}/2$ is nonzero. Using Eq. (3) and definition of $M$ in Eq. (19) we obtain Eq. (23).
We can also calculate the divergence of the full energy-momentum tensor in Eq. (23) using Eqs. (16) and (19) to find
@@
where $M^{\lambda\mu}=MB^{\lambda\mu}/B_{*}$ is the magnetization density due to the fast transverse motion of the particles. The current $J^{\mu}$ in Eq. (25) is the sum of the “transport current” $N^{\mu}$ and the “magnetization current” $\partial_{\lambda}M^{\lambda\mu}$. Substitution of $N_{\text{D}}$ from Eq. (21) converts Eq. (25) into Eq. (11b), i.e., only two out of four equations remain independent.
Constitutitive relations.—The conservation laws (11) form a system of three independent equations of guiding-center hydrodynamics. All components of $\tilde{N}^{\mu}$ and $T^{\mu}_{~{}\nu}$, thus, must be expressed through three independent variables via the constitutive relations.
At the ideal hydrodynamic level, the constitutive relations are obtained by substituting the equilibrium distribution functions $f_{\text{eq}}$ into the equation for the currents. This function $f_{\text{eq}}$ satisfies $\mathcal{C}[f_{\rm eq}]=0$ via detailed balance, which requires it to be the Fermi-Dirac or Bose-Einstein distribution function $f_{\rm eq}(g)=(e^{g}\mp 1)^{-1}$ with $g$ being a linear combination of the additive quantum numbers conserved in collisions, i.e., of energy-momentum and charge: $g=\beta u\cdot p-\alpha$. The coefficients $\alpha$ and $\beta u^{\mu}$ are related to the temperature and chemical potential via $T=1/\beta$ and $\mu=\alpha/\beta$. The unit 4-vector $u^{\mu}$ can be identified with the frame in which the total momentum vanishes. Using Eqs. (20) and (18), performing calculation in the frame $u=(1,\bm{0})$, where $g=\beta p_{0}-\alpha$, $p_{0}=\sqrt{p_{z}^{2}+\tilde{m}^{2}}$, and then boosting along the direction of the magnetic field, one obtains Eq. (13) with $n$, $\epsilon$, $P$, and $P_{\perp}$ given by the corresponding quantities in a one-dimensional gas, multiplied by the Landau-level density:
@@
Second-law of thermodynamics.—Ideal hydrodynamics must conserve entropy. We now show, without relying on kinetic description, i.e., without Eq. (26), that this requirement implies relationships (14) between the coefficients $n$, $\epsilon$, $P$ and $P_{\perp}$.
@@
Substituting Eqs. (13) into (27) and demanding that terms proportional to $\partial\!\cdot\!u$ vanish independently from terms with $u\cdot\partial$ derivatives, we find two relations:
@@
$ds+\alpha dn-\beta d\epsilon-\beta(P-P_{\perp})d\ln B_{*}=0,$ (28b)
where we have used $u^{\lambda}\partial_{\mu}\Delta^{\mu}_{\lambda}=-u^{\lambda}\partial_{\mu}\tilde{\Delta}^{\mu}_{\lambda}=(u\cdot\partial)\ln B_{*}$. Differentiating Eq. (28a) and comparing the result with Eq. (28b), one obtains Eqs. (14). In the kinetic theory, all relationships can be directly verified by taking derivatives of $P$ in Eq. (26).
Beyond kinetic theory.—Provided the frame velocity $u^{\mu}$ is defined via $B_{\mu\nu}u^{\nu}=0$, $\tilde{\Delta}^{\mu}_{\lambda}T^{\lambda}_{~{}\nu}u^{\nu}=\epsilon u^{\mu}$, the constitutive relations (13) are the most general covariant relations at zeroth order in derivatives. Therefore, the ideal hydrodynamic equations (11), (13) and second law constraints (14) are valid more generally, i.e., even when kinetic theory is not applicable, e.g., in the magnetized strongly coupled quark-gluon plasma.
@@
The constraint $B_{\mu\nu}u^{\nu}=0$ comes from finite conductivity $\sigma$ of the plasma, which damps motion in directions perpendicular to the magnetic field at the rate $\Gamma_{\perp}=\sigma B^{2}/(\epsilon+P)$, which is finite even at zero wave number. For small $B$ there is an intermediate range of time scales between the mean free time and $1/\Gamma_{\perp}$ where an extended hydrodynamic description with an unconstrained fluid velocity $u^{\mu}$ [22] is possible. This is an example of the Hydro+ regime [23], where the components of velocity perpendicular to the magnetic field play the role of additional non-hydrodynamic, but slow, modes. In the true hydrodynamic regime $u^{\mu}$ becomes constrained.
@@
The fact that there are two, not four, conservation laws for energy-momentum is a consequence of a symmetry: a crossed field $B_{\mu\nu}$ is still invariant under a two-parameter family of diffeomorphisms $x^{\mu}\to x^{\mu}+\xi^{\mu}$, such that $\xi^{\mu}B_{\mu\nu}=0$. This symmetry implies the conservation equations (11b). One can also derive the constitutive relations (13) from the partition function of the system in external metric and gauge fields [17].
@@
Conclusion.—In this Letter we have constructed a manifestly Lorentz covariant description of the guiding center motion based on an action principle. This allows one to formulate the kinetic theory and write down the ideal hydrodynamic equations of a fluid of charged particles in a background of high magnetic field. We expect our equation to be especially useful for relativistic particles with quantized transverse motion.

### v2

# Kinetic Theory and Hydrodynamics

@@ Abstract
We treat the guiding-center dynamics in a varying external Maxwell field using a relativistically covariant action principle which reproduces the known Vandervoort expression for the drift velocity and extends it to curved spacetime. We derive the corresponding kinetic theory and ideal hydrodynamic theory. In contrast to conventional five-equation hydrodynamics, the guiding-center hydrodynamics needs only three equations due to a constraint on the motion across magnetic field. We argue that such a hydrodynamics is applicable to strongly coupled plasmas where kinetic theory fails.
Introduction—About 50 years ago, Vitaly Ginzburg identified matter in ultrahigh magnetic fields as one of the most important and interesting problems in physics [1]. This problem remains current to this day. In astrophysics, magnetic fields exceeding $10^{15}~{}\text{G}$ have been found on the surface of a type of neutron star called a magnetar [2, 3]. In peripheral ultrarelativistic heavy-ion collisions, the magnetic field created by the colliding nuclei may exceed $10^{19}~{}\text{G}$ [4, 5] and leave measurable effects on the collisions [6]. In condensed matter physics, magnetic fields available in the lab ($\sim\!10^{5}~{}\text{G}$) can already be considered strong at low densities of charge carriers and temperatures, and lead to nontrivial phenomena like the integer and fractional quantum Hall effects [7, 8].
It is known that properties of matter change dramatically in high magnetic fields [9]. Ordinary atoms become deformed when the magnetic field $B$ reaches $m_{e}^{2}ce^{3}/\hbar^{3}\approx 2.35\times 10^{9}~{}\text{G}$ and, when $B$ exceeds $m_{e}^{2}c^{3}/(e\hbar)\approx 4.41\times 10^{13}~{}\text{G}$, the cyclotron motion of electrons is relativistic. Electron motion at high $B$ can be decomposed into a fast cyclotron motion and a slower motion of the orbit’s guiding center. At very high $B$, the cyclotron motion is quantum and, in the most extreme case, frozen onto the lowest Landau level. The motion of the guiding center, on the other hand, is (semi)classical and is a textbook problem of plasma physics. It is well known that the guiding center moves mostly along the magnetic field line with a slow “drift” off that line when the field is inhomogeneous [10]. This drift motion is important in various physical problems including cosmic-ray transport [11] and thermonuclear plasma confinement [12]. A plasma of electrons in a high $B$ field can be thought of as a plasma of guiding centers, and finding a description of the collective dynamics of these guiding centers is crucial for understanding the behavior of matter in very high magnetic fields. Guiding-center dynamics is used in simulations of astrophysical plasmas (see, e.g., Refs. [13, 14, 15, 16, 17]).
@@
Relativistic motion of the guiding center—In the special relativistic setting, the problem of guiding-center motion has been solved since the 1960s [18, 19] and revisited recently in Refs. [20, 21, 22, 23]. Here we provide an alternative, simpler, and manifestly Lorentz covariant description of the motion based on an action principle. This formulation is easy to generalize to curved spacetime and will turn out to be particularly apt for our subsequent derivation of the kinetic and hydrodynamic theories.
We assume the presence of a strong electromagnetic field $F_{\mu\nu}$ which is magnetically dominated, i.e., $F_{\mu\nu}F^{\mu\nu}=2(\bm{B}^{2}-\bm{E}^{2})>0$. At any given spacetime point there exists a frame where $\bm{E}\!\parallel\!\bm{B}$. We denote by $B_{*}$ and ${E_{*}}$ the magnitudes of the electric and magnetic fields in that frame (${E_{*}}<0$ if the fields are antiparallel). In terms of the Lorentz invariants, $F_{\mu\nu}F^{\mu\nu}=2(B_{*}^{2}-E_{*}^{2})$, $F_{\mu\nu}\tilde{F}^{\mu\nu}=-4E_{*}B_{*}$  [^1] We use the convention of Refs. [10, 45]: The metric signature is (+,,,)(+,-,-,-)( + , - , - , - ), $A^{\mu}=(\phi,\mathbf{A})$, $\tilde{F}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\alpha\beta}F_{\alpha\beta}$, where $\epsilon^{\mu\nu\alpha\beta}$ is the Levi-Civita tensor, $\epsilon^{0123}=+1$..
We consider the motion of an electron in this field and, to simplify further order-of-magnitude estimates, assume that the motion is relativistic: $v\sim 1$. Let $p_{\perp}$ be the magnitude of the electron’s momentum in the plane perpendicular the magnetic field. Then, the cyclotron radius is $a=p_{\perp}/(eB_{*})$. We assume $a$ is much smaller than the scale $L$ on which the magnetic field varies: $a\ll L$. In that case, the motion of the electron is a superposition of a fast cyclotron motion and the smoother motion of the guiding center. We can then expand in the gradients of the magnetic field, with each additional gradient being suppressed by $\mathcal{O}(a/L)$.
@@
$B_{\mu\nu}=\frac{F_{\mu\nu}-\epsilon_{*}\tilde{F}_{\mu\nu}}{1+\epsilon_{*}^{2}},\quad E_{\mu\nu}=\epsilon_{*}\frac{\tilde{F}_{\mu\nu}+\epsilon_{*}F_{\mu\nu}}{1+\epsilon_{*}^{2}},$ (1)
@@
The equations governing the motion of the guiding center were first derived in Ref. [18] (see also Ref. [19]) in explicit component form where Lorentz invariance is hidden. We show that these equations can be obtained straightforwardly from the manifestly Lorentz invariant least action principle for a relativistic particle with an additional constraint. Namely, let the trajectory of the particle in phase space be $(x^{\mu}(\tau),p_{\mu}(\tau))$, where $\tau$ is an affine parameter along the worldline. Consider the following Lagrangian:
$L=-\dot{x}\cdot(p+A)+\frac{\alpha}{2}(p^{2}-\tilde{m}^{2})+\lambda_{\mu}B^{\mu\nu}p_{\nu},$ (2)
@@
together with two constraints $p^{2}=\tilde{m}^{2}$ and $B^{\mu\nu}p_{\nu}=0$. From Eq. (5a) one sees that $\alpha$ and $\lambda_{\alpha}$ depend on the choice of the affine parameter $\tau$. For example, if $B_{\mu\nu}$ is purely spatial, choosing $\tau=t$ means that $\alpha=1/p_{0}$.
While the momentum of the guiding center always points in the direction of the magnetic field due to the constraint $B^{\mu\nu}p_{\nu}=0$, the particle does not move strictly along the magnetic field lines: It also slowly drifts in an orthogonal direction with a velocity given by the spatial components of $v_{\mathrm{D}}^{\mu}\equiv\lambda_{\alpha}B^{\alpha\mu}\,$. Substituting Eq. (5a) into Eq. (5b) and multiplying by $B^{\lambda\mu}$, we find
@@
where we dropped terms higher order in gradients of the magnetic field. Note that, since $v_{\mathrm{D}}$, and thus $\lambda$, is first order in gradients of $B$, the last term in Eq. (5b) is second order in the gradients.
Equation (6) compactly and covariantly summarizes the results of Refs. [18, 19]. The first and the second terms reproduce the drifts due to the change of the magnetic field direction (curvature drift) and magnitude (gradient drift), respectively. Additional effects due to time derivatives of the fields, dictated by Lorentz invariance, are also included [26].
The covariant generalization of Eqs. (2), (5), and (6) to curved spacetime is straightforward; see Eqs. (A1)–(A3) in the Appendix. In particular, the spacetime derivatives in Eq. (6) are simply replaced by covariant derivatives.
Kinetic equation—Using the equation of motion for a single particle derived above, we now study the kinetic theory of a collection of charged particles moving in a large external, almost crossed ($\epsilon_{*}\ll 1$), field. To be able to replace the particle motion by that of the guiding centers, the cyclotron radius must be shorter than the mean free path. For example, for a relativistic electron-positron plasma with $T\gtrsim m_{e}$, the cyclotron radius is $T/(eB)$, while the mean free path is (up to a log) $1/(e^{4}T)$. The kinetic theory we are going to derive works when $B\gg e^{3}T^{2}$.
@@
where the Liouville operator $\mathcal{L}[f]$ represents the rate of change (per unit worldline “time”) of the occupation number $f(x,p)$ along the particle worldline determined by Eqs. (5). The occupation number is changed by collisions; therefore, the rate is equal to the collision integral $\mathcal{C}[f]$.
To describe transport in the gas, we need to know also the invariant phase-space integration measure, i.e., the number of single-particle states in a given phase-space volume at a given time in the lab frame. The number of states is time independent if the phase-space volume evolves according to the equations of motion. This statement can be expressed in Lorentz covariant form by introducing the conserved phase-space 8-current $J^{M}$ given by [27]
@@
i.e., by the product of the phase-space “velocity” and the “density” $W$  [^3] We need to extend the velocities $\dot{x}^{\mu}$, $\dot{p}_{\nu}$ to the whole eight-dimensional phase space. We can do that by simply declaring Eqs. (5) and (6) to be valid everywhere, even outside the constraint hypersurface.
@@
$W=\frac{B_{*}+B_{*}^{-1}p_{\mu}\partial_{\nu}B^{\mu\nu}}{2\pi^{2}\alpha}\theta(p_{0})\delta(p^{2}-\tilde{m}^{2})\delta^{2}(\Delta^{\mu\nu}p_{\nu}),$ (10)
where we have defined the projector $\Delta^{\lambda}_{\nu}$ as
For a magnetic field along the $z$ axis, $\Delta^{\mu}_{\nu}$ projects onto the $(x,y)$ plane, while the complement $\tilde{\Delta}^{\mu}_{\nu}$, onto the $(t,z)$ hyperplane.
Note that the equation $\partial_{M}J^{M}=0$ determines $W$ only up to an overall constant. This constant is chosen in Eq. (10) to match the Landau-level degeneracy.
Hydrodynamics—We can now use the guiding-center kinetic theory to describe the transport in the hydrodynamic regime. For simplicity, we consider a single-component plasma neutralized by an inert background charge density. The electromagnetic field is treated mostly as a fixed background, but can be made dynamical by supplementing the hydrodynamic equations with Maxwell’s equations.
@@
The presence of the projector $\tilde{\Delta}$ means that Eq. (12b) represents only two independent equations, not four. These conservation laws are supplemented by the constitutive relations that express current $\tilde{N}^{\mu}$ and stress tensor $T^{\mu\nu}$ in terms of three independent hydrodynamic variables: the local temperature $T(x)$, local chemical potential $\mu(x)$, and local velocity $u^{\mu}(x)$. The velocity is subject to the constraints $u\cdot u=1$, $B_{\mu\nu}(x)u^{\nu}(x)=0$ and hence represents only 1 degree of freedom.
@@
The anisotropy of the pressure tensor due to the magnetic field seen in Eq.(13b) was discussed, e.g., in Refs. [29, 30, 31, 32, 33].
Derivation of hydrodynamics—Multiplying Eq. (7) by $W$ and $Wp_{\nu}$ and integrating over $p$, using the fact that the collision term $\mathcal{C}[f]$ conserves the particle number (charge) and 4-momentum, i.e.  [^4] We have chosen the charge of particles to be 1. If there is more than one type of particle, one needs to sum over particle types, labeled by $a$. For example, conservation of charge becomes $\sum_{a}\int_{p}Wq_{a}C[f_{a}]=0$ where $q_{a}$ is the charge of the particles of type $a$. Similarly, if the values of adiabatic invariant $J$ vary, they have to be summed over too, e.g., in the second equation of (19).,
@@
$F_{\nu}=F_{\nu\lambda}N^{\lambda}-M\partial_{\nu}B_{*},~{}~{}M\equiv-\!\int_{p}W\frac{\alpha J}{2}f$ (19)
@@
We will call $N^{\mu}$ in Eq. (17) the “transport current” to distinguish it from the full current, to be introduced later. It consists of two parts: $N^{\mu}=\tilde{N}^{\mu}+N_{\text{D}}^{\mu}$, where the longitudinal current
$\tilde{N}^{\mu}\equiv\int_{p}W\alpha p^{\mu}f$ (20)
flows along the magnetic field lines ($B_{\mu\nu}\tilde{N}^{\nu}=0$), while the drift current
$N_{\text{D}}^{\mu}\equiv\int_{p}Wv_{\text{D}}^{\mu}f=B_{*}^{-2}(\tilde{T}^{\lambda}_{\nu}\partial_{\lambda}B^{\mu\nu}-MB^{\mu\nu}\partial_{\nu}B_{*})$ (21)
@@
Then, Eq. (22) takes a simper form of Eq. (12b[35].
Physically, the contribution $T_{\perp}^{\mu\nu}$ in Eq. (23) comes from the fast transverse motion of the particles absent in Eq. (18). Indeed,
@@
While $\overline{p_{\perp}^{\mu}}=0$, $\overline{p_{\perp}^{\mu}p_{\perp}^{\nu}}=-\Delta^{\mu\nu}p_{\perp}^{2}/2$ is nonzero. Using Eq. (3) and the definition of $M$ in Eq. (19) we obtain Eq. (23).
We can also calculate the divergence of the full energy-momentum tensor in Eq. (23) using Eqs. (16) and (19):
@@
where $M^{\lambda\mu}=MB^{\lambda\mu}/B_{*}$ is the magnetization density due to the fast transverse motion of the particles. The full current $J^{\mu}$ in Eq. (25) is the sum of the transport current $N^{\mu}=\tilde{N}^{\mu}+N_{\text{D}}^{\mu}$ and the “magnetization current” $\partial_{\lambda}M^{\lambda\mu}$. The substitution of $N_{\text{D}}$ from Eq. (21) converts Eq. (25) into Eq. (12b); i.e., only two out of four equations remain independent.
Constitutive relations—The conservation laws (12) form a system of three independent equations of guiding-center hydrodynamics. All components of $\tilde{N}^{\mu}$ and $T^{\mu}_{~{}\nu}$, thus, must be expressed through three independent variables via the constitutive relations.
At the ideal hydrodynamic level, the constitutive relations are obtained by substituting the equilibrium distribution functions $f_{\text{eq}}$ into the equation for the currents. This function $f_{\text{eq}}$ satisfies $\mathcal{C}[f_{\rm eq}]=0$ via detailed balance, which requires it to be the Fermi-Dirac or Bose-Einstein distribution function $f_{\rm eq}(g)=(e^{g}\mp 1)^{-1}$ with $g$ being a linear combination of the additive quantum numbers conserved in collisions, i.e., of 4-momentum and charge: $g=\beta u\cdot p-\alpha$. The coefficients $\alpha$ and $\beta u^{\mu}$ are related to the temperature and chemical potential via $T=1/\beta$ and $\mu=\alpha/\beta$. The unit 4-vector $u^{\mu}$ can be identified with the frame in which the total 3-momentum vanishes. Using Eqs. (20) and (18), and performing a calculation in the frame $u=(1,\bm{0})$, where $g=\beta p_{0}-\alpha$, $p_{0}=\sqrt{p_{z}^{2}+\tilde{m}^{2}}$, and then boosting along the direction of the magnetic field, one obtains Eq. (13) with $n$, $\epsilon$, $P$, and $P_{\perp}$ given by the corresponding quantities in a one-dimensional gas, multiplied by the Landau-level density:
@@
Second law of thermodynamics—Ideal hydrodynamics must conserve entropy. We now show, without relying on kinetic description, i.e., without Eq. (26), that this requirement implies relationships (14) between the coefficients $n$, $\epsilon$, $P$ and $P_{\perp}$.
@@
Substituting Eqs. (13) into (27) and demanding that terms proportional to $\partial\!\cdot\!u$ vanish independently from terms with $u\cdot\partial$ derivatives, we find two relations [35]:
@@
$ds+\alpha dn-\beta d\epsilon-\beta(P-P_{\perp})d\ln B_{*}=0.$ (28b)
Differentiating Eq. (28a) and comparing the result with Eq. (28b), one obtains Eqs. (14). In the kinetic theory, all relationships can be directly verified by taking derivatives of $P$ in Eq. (26).
Beyond kinetic theory—Provided the frame velocity $u^{\mu}$ is defined via $B_{\mu\nu}u^{\nu}=0$, $\tilde{\Delta}^{\mu}_{\lambda}T^{\lambda}_{~{}\nu}u^{\nu}=\epsilon u^{\mu}$, the constitutive relations (13) are the most general covariant relations at zeroth order in derivatives. Therefore, the ideal hydrodynamic equations (12) and (13) and second law constraints (14) could be valid in the magnetized strongly coupled quark-gluon plasma, where kinetic theory fails.
@@
The constraint $B_{\mu\nu}u^{\nu}=0$ comes from finite conductivity $\sigma$ of the plasma, which damps motion in directions perpendicular to the magnetic field at the rate $\Gamma_{\perp}=\sigma B^{2}/(\epsilon+P)$, which is finite even at zero wave number. For small $B$ there is an intermediate range of timescales between the mean free time and $1/\Gamma_{\perp}$ where an extended hydrodynamic description with an unconstrained fluid velocity $u^{\mu}$ [31] is possible. This is an example of the Hydro+ regime [36], where the components of velocity perpendicular to the magnetic field play the role of additional nonhydrodynamic, but slow, modes. In the true hydrodynamic regime, $u^{\mu}$ becomes constrained.
@@
The fact that there are two, not four, conservation laws for 4-momentum is a consequence of a symmetry: A crossed field $B_{\mu\nu}$ is still invariant under a two-parameter family of diffeomorphisms $x^{\mu}\to x^{\mu}+\xi^{\mu}$, such that $\xi^{\mu}B_{\mu\nu}=0$. This symmetry implies the conservation equations (12b). One can also derive the constitutive relations (13) from the partition function of the system in external metric and gauge fields [26].
@@
Conclusion—In this Letter we constructed a manifestly relativistically covariant description of the guiding-center motion based on an action principle. This allows one to formulate the kinetic theory and write down the ideal hydrodynamic equations of a fluid of charged particles in a background of high magnetic field. We expect our equation to be especially useful for relativistic particles with quantized transverse motion.